OpenStudy (anonymous):

The derivative of x/(2 - tan x) = [-tan x + xsec(^2)x + 2]/(2 - tan x)^2 OR /(tan x - 2)^2 ????

OpenStudy (phi):

in general, (a-b)^2 = (b-a)^2 because a-b= -(a-b) (i.e. -a+b or b-a) when you square a or -a you get a^2 so the sign does not matter.